# backgroundcolor=gray!3, roundcorner=3pt

### v1

@@ Abstract
Human guidance is often desired in reinforcement learning to improve the performance of the learning agent. However, human insights are often mere opinions and educated guesses rather than well-formulated arguments. While opinions are subject to uncertainty, e.g., due to partial informedness or ignorance about a problem, they also emerge earlier than hard evidence could be produced. Thus, guiding reinforcement learning agents through opinions offers the potential for more performant learning processes, but comes with the challenge of modeling and managing opinions in a formal way. In this article, we present a method to guide reinforcement learning agents through opinions. To this end, we provide an end-to-end method to model and manage advisors’ opinions. To assess the utility of the approach, we evaluate it with synthetic and human advisors, at different levels of uncertainty, and under multiple advise strategies. Our results indicate that opinions, even if uncertain, improve the performance of reinforcement learning agents, resulting in higher rewards, more efficient exploration, and a better reinforced policy. Although we demonstrate our approach in a simplified topological running example, our approach is applicable to complex problems with higher dimensions as well.
@@ 1 Introduction
General guidance-based techniques allow humans to inform agents about future aspects of the task, e.g., interesting regions to explore [47] or a trajectory (i.e., a sequence of actions) to follow [50]. A severe limitation of such techniques is the inconsistent correctness of human advice. As Scherf et al [45] report, the majority of interactive or guided reinforcement learning approaches assume the human input to be useful and correct [36]. However, this is not always true in real applications [31, 34, 4]. Informed advisors, such as experts, often express advice in the form of opinions—cognitive constructs that are subject to epistemic uncertainty, i.e., uncertainty that is an artifact of a lack of knowledge, partial informedness, and ignorance. Opinions about the solutions or constraints of a problem emerge earlier than hard evidence can be produced. Thus, relying on opinions to provide advice to reinforcement learning agents results in more agile and efficient reinforcement learning methods. For this, epistemic uncertainty has to be approached in a formal way in order to avoid unsound advice that would adversely affect the agent. Unfortunately, despite the fact that the importance of dealing with uncertainty in machine learning has been well-recognized [26], there is a lack of reinforcement learning methods that treat epistemic uncertainty as a first-class citizen in guided reinforcement learning. Specifically, methods for managing epistemic uncertainty are missing [26].
@@
Subjective logic [28] is an extension of probabilistic logic [1], in which users can express subjective opinions by quantified parameters of belief and certainty. Opinions are formed from a belief component and an uncertainty component. In statistics and economics, the uncertainty component of subjective logic is often called second-order probability [22], and is represented in terms of a probability density function over first-order probabilities. Unfortunately, the complex framework of subjective logic limits its applicability. Most techniques that rely on it resort to simplifications, e.g., assuming that humans can express their opinions in terms of mathematical abstractions [10]; or emulating human opinion by objectively measured metrics, taking away the core subjective element of the approach [52].
Goal and research questions. Our goal is to evaluate the feasibility and utility of guiding RL agents by opinions—i.e., through advice that is subject to belief uncertainty—captured in constructs of subjective logic. To meet our goal, we formulate the following research questions.
@@
We develop a method in which opinion is formulated through the sound mathematical foundations of subjective logic, and fused into the policy of reinforcement learning agents. Subsequently, we evaluate the performance of the advised agent through a series of experiments. Our results indicate that opinions, even if uncertain, improve the performance of reinforcement learning agents, resulting in higher rewards, more efficient exploration, and a better reinforced policy. Since opinions emerge earlier and more easily than hard evidence could be produced, opinion-guided reinforcement learning offers a more efficient and economical alternative to traditional guided reinforcement learning approaches.
Contributions. The main contribution of this work is a modeling approach for guiding reinforcement learning through subjective opinions, i.e., advice that is subject to belief uncertainty. We investigate various flavors of reinforcement learning algorithms and show where and how to augment them with hints expressed via subjective logic. Since the human-in-the-loop can quickly become a bottleneck in human-machine collaboration due to its expensive reward function [13], we also recommend architectural patterns to implement opinion-guided RL. Through detailed experiments, we show that the performance of RL agents improves with added human input and improves again with added uncertainty and disbelief information. Our approach fosters more efficient collaboration between human and machine agents [37] while improving the performance of RL agents.
Structure. The rest of this paper is structured as follows. In Sec. 2, we present the running example we use throughout the paper to illustrate key concepts, and to drive the evaluation of our approach. In Sec. 3, we briefly review the background topics relevant to our work and review the related work. In Sec. 4, we outline our approach to modeling human belief and certainty in reinforcement learning. In Sec. 5, we evaluate our approach under various advice types, sources, and learning parameters. In Sec. 6, we discuss the results and their implications and outline a set of open challenges. In Sec. 7, we draw the conclusions and outline future work.
@@ 2 Running example
Stepping on a frozen tile Nothing happens; the agent is allowed to continue the traversal of the environment. 0
Stepping in a hole The task ends unsuccessfully. -1
@@
The goal of the agent is to learn the optimal path from the top leftmost tile to the bottom rightmost tile under the given reward structure. Through trial-and-error, the agent gradually learns which actions are beneficial in a specific situation. For example, in cell (2, 2), the agent should not move to the right, because this action would result in stepping into a hole and terminating the exploration unsuccessfully, with a reward of -1. Instead, in cell (2, 2), the agent should move down, because cell (3, 2) is a frozen tile, and it it allows the agent to move closer to the goal, which is in cell (3, 3).
@@ 3.1 Reinforcement learning
3.1.1 Various flavors of reinforcement learning
While all RL algorithms aim maximize expected return, various methods achieve this goal in different ways. Value-based methods aim to maximize either the state-value function $v_{\pi}(s)=\mathbb{E}_{\pi}[G_{t}|s]$, the action-value function $q_{\pi}(s,a)=\mathbb{E}_{\pi}[G_{t}|s,a]$, or both. Value-based approaches are deterministic because they select actions greedily when maximizing the value function; however, this might lead to under-exploration. In contrast, the goal of policy-based methods is to learn a parameterized policy that maximizes the reward function in every state, without the use of value functions. Policy-based methods explore the state space more thoroughly than value-based methods, but produce estimates with more noise, potentially resulting in unstable learning processes. Finally, actor-critic methods provide a trade-off between value-based and policy-based methods, in which the actor implements a policy-based strategy, and the critic criticizes the actions made by the actor based on a value function. Different methods offer various benefits and perform with unique utility in specific learning problems.
In particular, policy gradient is a subset of the aforementioned policy-based methods. In policy gradient, the policy is parameterized through a parameter vector $\boldsymbol{\theta}$. For reasonably sized action/state spaces, the policy may be parameterized as numerical preferences for all state-action pairs $h(s,a,\boldsymbol{\theta})\in\mathbb{R}$, and actions can then be chosen using a soft-max distribution. This is called discrete policy gradient, and often, a matrix-like representation of a look-up table is used. In contrast, the policy may be instead parameterized via a neural network such that $\boldsymbol{\theta}$ is a vector of connection weights of the network, or linear in features, which defines deep policy gradient. In either case, the policy is defined as $\pi(a|s,\boldsymbol{\theta})$, the probability of taking action $a$ in state $s$ with parameter $\boldsymbol{\theta}$. The policy may be parameterized in any manner, subject to the condition that it is differentiable with respect to $\boldsymbol{\theta}$. This way, $\boldsymbol{\theta}$ may be updated to maximize some scalar performance metric $J(\boldsymbol{\theta})$.
@@ 3.2 Subjective logic
To formalize human guidance in reinforcement learning, we rely on subjective logic. Subjective logic [28] is an extension of probabilistic logic [1], in which users can express subjective opinions by quantified parameters of belief and certainty. Quantifying subjective opinions is an improvement over traditional logic systems as it promotes opinions to first-class citizens in a systematic fashion. Well-informed opinions can be early indicators of emerging new knowledge and often, opinions are the only available insights in engineering. Subjective logic has seen growing applications, such as in model-driven software engineering, e.g., by Burgueño et al [10], where it is used to model belief uncertainty [5] in domain models. As well, it has been used in knowledge graphs as per Navarrete and Vallecillo [42] to allow reasoning about graph databases enriched with uncertainty.
@@ 3.2.1 Formal underpinnings
Given a boolean predicate $x$, a binomial opinion regarding the truth of $x$ is given by $\omega_{x}=(b_{x},d_{x},u_{x},a_{x})$, where $b_{x}$ represents the belief in the truthfulness of $x$; $d_{x}$ represents the disbelief in the truthfulness of $x$; $u_{x}$ quantifies the degree of epistemic uncertainty or uncommitted belief concerning $x$; and $a_{x}$ represents the base rate, i.e., the prior probability of $x$ in the absence of (dis)belief. For each parameter, $0\leq b_{x},d_{x},u_{x},a_{x}\leq 1$. These parameters satisfy the following conditions.
@@
Equation 2 expresses the projected probability of an opinion, effectively transforming subjective opinions into the probability domain, where $P(x)$ means the probability that boolean predicate $x$ holds. It is clear that as uncertainty $u_{x}$ increases, projected probability $P_{x}$ is closer to base rate $a_{x}$. In contrast, as uncertainty $u_{x}$ decreases, projected probability $P_{x}$ is closer to that of the belief parameter $b_{x}$. Both Equations 1 and 2 are important invariants that we rely on throughout the paper, particularly in Sections 4 and 5.
Informed opinions work well at scale. That is, the more experts that express their opinions, the higher the credibility of a collective opinion. Collective opinions can be inferred by semantically sound fusion operators [30]. A fusion operator is a function $f:\Omega\times\Omega\to\Omega$ that maps a pair of subjective opinions onto a new, joint subjective opinion. Various fusion semantics exist, and the right operator must be chosen based on its fit for purpose. A detailed account of fusion operators is given in Sec. 4.3.3.
@@ 3.2.2 Challenges in employing subjective logic
The high expressive power of subjective logic comes with the added challenge of calibrating its parameters, particularly uncertainty $u$ and base rate $a$. There are two schools of thought to address this problem. In permissive techniques, setting uncertainty and belief-disbelief are deferred to the user [10, 27], and it is assumed that these parameters will become available at some point. However, it is usually not explained when, how, and who has to set these parameters, leading to a limited applicability of subjective logic. In restrictive techniques, users are not asked to parameterize the framework. Instead, parameters are controlled by external measures [29, 39]. Unfortunately, this takes away subjective elements from subjective logic. Margoni and Walkinshaw [39] and Walkinshaw and Shepperd [52] use statistical evidence to set values, employing subjective logic to analyze empirical studies and experimental results, respectively. In regards to base rate $a_{x}$, some sources opt to set this value via statistical evidence, such as in Jøsang and Bondi [29], Margoni and Walkinshaw [39]. Navarrete and Vallecillo [42] allow users to directly indicate the base rate, and Walkinshaw and Shepperd [52] assume $a_{x}=0.5$.
@@ 3.3 Related work
The closest work to ours is the one by Guo et al [24], who combine feedback from several human trainers to create a reliable reward for the RL agent. SL is used to model the agent’s opinion or trust of the human trainer. Converting this opinion into a probability yields a measure of the trainer’s trustworthiness, and after the agent utilizes human feedback, the agent’s opinion about the trainer’s is updated. In contrast, our work uses SL the other way around: to model the advisor’s opinion about the environment, and uses that opinion to drive policy shaping and improve the agent’s performance.
Other related works focus either on guiding reinforcement learning agents but without the use of subjective logic (Sec. 3.3.2) and employing subjective logic in reinforcement learning but not for guidance (Sec. 3.3.1).
@@ 3.3.1 Subjective logic in reinforcement learning (but not for guidance)
The combination of RL and SL is present in a number of publications. Notable work includes Zennaro and Jøsang [56], an initial inquiry into using SL in multi-armed bandits, a simplified RL problem. In this work, the agent forms an opinion over the available actions, and if the agent receives a positive reward for an action, the opinion about said action can be updated. Subjective logic to capture uncertainty in multi-armed bandit like problems is extended in Wang and Grace [53]. Similarly, Guo et al [25] uses a Deep Reinforcement Learning (DRL) multi-agent in smart farm monitoring to collect data from sensors on cattle- opinions about each animal’s condition is computed via this sensor data. In [58], RL is used to learn a low dimensional representation of safe regions of complex dynamical systems. Safety of states is estimated using DSAF (Discretized Safety Assessment Function), which is described with SL. In Guo et al [25], SL is used to model and compute uncertainty in a DRL framework used to identify the intents of tweets. In Zhao et al [57], opinions are used to represent edges in a graph network, and for unknown edges, a set of best paths is determined via DRL. This work proposes 3 different DRL models, each with reward given based on different types of uncertainty that are computed via SL. Similarly, Zhao et al [57] uses SL to represent users within an Online Social Network. Different types of users are initialized with particular opinions, which can be updated via SL fusion operators to enhance DRL based Competitive Influence Maximization. While these works combine RL and SL, our work differs by modeling both RL policy and user advice in SL, and using this model to guide the SL agent via human opinion.
3.3.2 Guidance in Reinforcement Learning (without SL)

### v2

@@ Abstract
Human guidance is often desired in reinforcement learning to improve the performance of the learning agent. However, human insights are often mere opinions and educated guesses rather than well-formulated arguments. While opinions are subject to uncertainty, e.g., due to partial informedness or ignorance about a problem, they also emerge earlier than hard evidence can be produced. Thus, guiding reinforcement learning agents by way of opinions offers the potential for more performant learning processes, but comes with the challenge of modeling and managing opinions in a formal way. In this article, we present a method to guide reinforcement learning agents through opinions. To this end, we provide an end-to-end method to model and manage advisors’ opinions. To assess the utility of the approach, we evaluate it with synthetic (oracle) and human advisors, at different levels of uncertainty, and under multiple advice strategies. Our results indicate that opinions, even if uncertain, improve the performance of reinforcement learning agents, resulting in higher rewards, more efficient exploration, and a better reinforced policy. Although we demonstrate our approach through a two-dimensional topological running example, our approach is applicable to complex problems with higher dimensions as well.
@@ 1 Introduction
General guidance-based techniques allow humans to inform agents about future aspects of the task, e.g., interesting regions to explore [53] or a trajectory (i.e., a sequence of actions) to follow [56]. A severe limitation of such techniques is the inconsistent correctness of human advice. As Scherf et al [51] report, the majority of interactive or guided RL approaches assume the human input to be useful and correct [41]. However, this is not always true in real applications [36, 39, 4]. Informed advisors, such as experts, often express advice in the form of opinions—cognitive constructs that are subject to epistemic uncertainty, i.e., uncertainty that is an artifact of a lack of knowledge, partial informedness, and ignorance. Opinions about the solutions or constraints of a problem emerge earlier than hard evidence can be produced. Thus, relying on opinions to provide advice to RL agents results in more agile and efficient RL methods. For this, epistemic uncertainty has to be approached in a formal way in order to avoid unsound advice that would adversely affect the agent. Unfortunately, despite the fact that the importance of dealing with uncertainty in machine learning has been well-recognized [30], there is a lack of RL methods that treat epistemic uncertainty as a first-class citizen in guided RL. Specifically, methods for managing epistemic uncertainty are missing [30].
@@
Subjective logic [33] is an extension of probabilistic logic [1], in which users can express opinions by quantified parameters of belief and certainty. Opinions are formed from a belief component and an uncertainty component. In statistics and economics, the uncertainty component of subjective logic is often called second-order probability [25], and is represented in terms of a probability density function over first-order probabilities. Unfortunately, the complex framework of subjective logic limits its applicability. Most techniques that rely on it resort to simplifications, e.g., assuming that humans can express their opinions in terms of mathematical abstractions [11]; or emulating human opinion by objectively measured metrics, taking away the core subjective element of the approach [58].
Goal and research questions
Our goal is to evaluate the feasibility and utility of guiding RL agents by opinions—i.e., through advice that is subject to belief uncertainty—captured in constructs of subjective logic. To meet our goal, we formulate the following research questions.
@@
We develop a method in which opinions are formulated through the sound mathematical foundations of subjective logic, and fused into the policy of RL agents. Subsequently, we evaluate the performance of the advised agent through a series of experiments. Our results indicate that opinions, even if uncertain, improve the performance of RL agents, resulting in higher rewards, more efficient exploration, and a better reinforced policy. Since opinions emerge earlier and more easily than hard evidence can be produced [18], opinion-guided RL offers a more efficient and economical alternative to traditional guided RL approaches.
Contributions
The main contribution of this work is a modeling approach for guiding RL through opinions, i.e., advice that is subject to belief uncertainty. We investigate various flavors of RL algorithms and show where and how to augment them with advice expressed via subjective logic. Since the human-in-the-loop can quickly become a bottleneck in human-machine collaboration due to its expensive reward function [14], we also recommend architectural patterns to implement opinion-guided RL. Through detailed experiments, we show that the performance of RL agents improves with added human input and improves again with added uncertainty and disbelief information. Our approach fosters more efficient collaboration between human and machine agents [42] while improving the performance of RL agents.
Structure
The rest of this paper is structured as follows. In Sec. 2, we present the running example we use throughout the paper to illustrate key concepts, and to drive the evaluation of our approach. In Sec. 3, we briefly review the background topics relevant to our work and review the related work. In Sec. 4, we outline our approach to modeling human belief and certainty in RL. In Sec. 5, we evaluate our approach under various advice types, sources, and learning parameters. In Sec. 6, we discuss the results and their implications and outline a set of open challenges. In Sec. 7, we reflect on the threats to the validity of this study and discuss the mitigation strategies we applied. In Sec. 8, we draw the conclusions and outline future work.
@@ 2 Running example
Encountering a frozen tile Nothing happens; the agent is allowed to continue the traversal of the environment. 0
Encountering a hole The task ends unsuccessfully. 0
@@
The goal of the agent is to learn the optimal path from the top leftmost tile to the bottom rightmost tile under the given reward structure. Through trial-and-error, the agent gradually learns which actions are beneficial in a specific situation. For example, in cell [2,2], the agent should not move to the right, because this action would result in stepping into a hole and terminating the exploration unsuccessfully, with a reward of 0. Instead, in cell [2,2], the agent should move down, because cell [3,2] is a frozen tile, and it it allows the agent to move closer to the goal, which is in cell [3,3].
@@ 3.1 Reinforcement learning
3.1.1 Various flavors of RL
While all RL algorithms aim to maximize expected return, various methods achieve this goal differently. Value-based methods aim to maximize either the state-value function $v_{\pi}(s)=\mathbb{E}_{\pi}[G_{t}|s]$, the action-value function $q_{\pi}(s,a)=\mathbb{E}_{\pi}[G_{t}|s,a]$, or both. Value-based approaches are deterministic because they select actions greedily when maximizing the value function, which might lead to under-exploration. In contrast, the goal of policy-based methods is to learn a parameterized policy that maximizes the reward function in every state without using value functions. Policy-based methods explore the state space more thoroughly than value-based methods, but produce estimates with more noise, potentially resulting in unstable learning processes. Finally, actor-critic methods provide a trade-off between value-based and policy-based methods, in which the actor implements a policy-based strategy, and the critic criticizes the actions made by the actor based on a value function. Different methods offer various benefits and perform with unique utility in specific learning problems.
In particular, policy gradient is a subset of the aforementioned policy-based methods. In policy gradient, the policy is parameterized through a parameter vector $\boldsymbol{\theta}$. For reasonably sized action/state spaces, the policy may be parameterized as numerical preferences for all state-action pairs $h(s,a,\boldsymbol{\theta})\in\mathbb{R}$, and actions can then be chosen using a soft-max distribution. This is called discrete policy gradient, and often, a matrix-like representation of a look-up table is used. In contrast, the policy may be parameterized via a neural network such that $\boldsymbol{\theta}$ is a vector of network connection weights or linear in-features, which defines deep policy gradient. In either case, the policy is defined as $\pi(a|s,\boldsymbol{\theta})$, the probability of taking action $a$ in state $s$ with parameter $\boldsymbol{\theta}$. The policy may be parameterized in any manner, subject to the condition that it is differentiable with respect to $\boldsymbol{\theta}$. This way, $\boldsymbol{\theta}$ may be updated to maximize some scalar performance metric $J(\boldsymbol{\theta})$.
@@ 3.2 Subjective logic
To formalize human guidance in RL, we rely on subjective logic. Subjective logic [33] is an extension of probabilistic logic [1], in which users can express opinions by quantified parameters of belief and certainty. Quantifying opinions is an improvement over traditional logic systems as it promotes opinions to first-class citizens in a systematic fashion. Well-informed opinions can be early indicators of emerging new knowledge and often, opinions are the only available insights in engineering. Subjective logic has seen growing applications, such as in model-driven software engineering, e.g., by Burgueño et al [11], where it is used to model belief uncertainty [6] in domain models. As well, it has been used in knowledge graphs as per Navarrete and Vallecillo [48] to allow reasoning about graph databases enriched with uncertainty.
@@ 3.2.1 Formal underpinnings
Given a boolean predicate $x$, a binomial opinion regarding the truth of $x$ is given by $\omega_{x}=(b_{x},d_{x},u_{x},a_{x})$, where $b_{x}$ represents the belief in the truthfulness of $x$; $d_{x}$ represents the disbelief in the truthfulness of $x$; $u_{x}$ quantifies the degree of epistemic uncertainty or uncommitted belief concerning the truthfulness of $x$; and $a_{x}$ represents the base rate, i.e., the prior probability of $x$ being true in the absence of (dis)belief. For each parameter, $0\leq b_{x},d_{x},u_{x},a_{x}\leq 1$. These parameters satisfy the following conditions.
@@
Equation 2 expresses the projected probability of an opinion, effectively transforming opinions into the probability domain, where $P(x)$ means the probability that boolean predicate $x$ holds. It is clear that as uncertainty $u_{x}$ increases, projected probability $P_{x}$ is closer to base rate $a_{x}$. In contrast, as uncertainty $u_{x}$ decreases, projected probability $P_{x}$ is closer to that of the belief parameter $b_{x}$. Both Equations 1 and 2 are important invariants that we rely on throughout the paper, particularly in Sections 4 and 5.
A traditional probability $p$ corresponds with $u_{x}=0$, and can thus be transformed into a binomial opinion as follows.
$\omega_{x}=(p,1-p,0,p)$ (3)
Informed opinions work well at scale. That is, the more experts that express their opinions, the higher the credibility of a collective opinion. Collective opinions can be inferred by semantically sound fusion operators [35]. A fusion operator is a function $f:\Omega\times\Omega\to\Omega$ that maps a pair of opinions onto a new, fused opinion. Various fusion semantics exist, and the right operator must be chosen based on its fit for purpose. A detailed account of fusion operators is given in Sec. 4.3.3.
@@ 3.2.2 Challenges in employing subjective logic
The high expressive power of subjective logic comes with the added challenge of calibrating its parameters, particularly uncertainty $u$ and base rate $a$. There are two schools of thought to address this problem. In permissive techniques, setting uncertainty and belief-disbelief are deferred to the user [11, 32], and it is assumed that these parameters will become available at some point. However, it is usually not explained when, how, and who has to set these parameters, leading to a limited applicability of subjective logic. In restrictive techniques, users are not asked to parameterize the framework. Instead, parameters are controlled by external measures [34, 44]. Unfortunately, this takes away subjective elements from subjective logic. Margoni and Walkinshaw [44] and Walkinshaw and Shepperd [58] use statistical evidence to set values, employing subjective logic to analyze empirical studies and experimental results, respectively. In regards to base rate $a_{x}$, authors often opt to set this value via statistical evidence [34, 44]. Navarrete and Vallecillo [48] allow users to directly indicate the base rate, and Walkinshaw and Shepperd [58] assume $a_{x}=0.5$.
@@ 3.3 Related work
The closest ours is the work fo Guo et al [28], who combine feedback from several human trainers to create a reliable reward for the RL agent. Subjective logic is used to model the agent’s opinion or trust of the human trainer. Converting this opinion into a probability yields a measure of the trainer’s trustworthiness, and after the agent utilizes human feedback, the agent’s opinion about the trainers is updated. In contrast, our work uses subjective logic the other way around: to model the advisor’s opinion about the environment, and uses that opinion to drive policy shaping and improve the agent’s performance.
Other related works focus either on guiding RL agents but without the use of subjective logic (Sec. 3.3.2) and employing subjective logic in RL but not for guidance (Sec. 3.3.1).
@@ 3.3.1 Subjective logic in reinforcement learning (but not for guidance)
The combination of RL and subjective logic is present in a number of publications. Notable work includes the one by Zennaro and Jøsang [62], an initial inquiry into using subjective logic in multi-armed bandits, a simplified RL problem. In this work, the agent forms an opinion over the available actions, and if the agent receives a positive reward for an action, the opinion about said action can be updated. Subjective logic to capture uncertainty in multi-armed bandit like problems is extended by Wang and Grace [59]. Similarly, Guo et al [29] use a Deep Reinforcement Learning (DRL) multi-agent in smart farm monitoring to collect data from sensors on cattle. Opinions about each animal’s condition are computed via this sensor data. [65] use RL to learn a low dimensional representation of safe regions of complex dynamical systems. Safety of states is estimated using DSAF (Discretized Safety Assessment Function), which is described with subjective logic. Guo et al [29] use subjective logic to model and compute uncertainty in a DRL framework used to identify the intents of tweets. Zhao et al [64] use opinions to represent edges in a graph network, and for unknown edges, a set of best paths is determined via DRL. This work proposes 3 different DRL models, each with reward given based on different types of uncertainty that are computed via subjective logic. Similarly, Zhao et al [64] use subjective logic to represent users within an Online Social Network. Different types of users are initialized with particular opinions, which can be updated via subjective logic fusion operators to enhance DRL based Competitive Influence Maximization. While these works combine RL and subjective logic, our work differs by modeling both RL policy and user advice in subjective logic, and using this model to guide the subjective logic agent via human opinion.
3.3.2 Guidance in reinforcement learning (without Subjective Logic)